OpenStudy (anonymous):

How many 4 person committes are possible from a group of 9 people if there are no restrictions?

OpenStudy (anonymous):

\[C_{4}^{10}\]..by combination

OpenStudy (anonymous):

ie.10!/4! (10-4)!=210..i guess so..

OpenStudy (anonymous):

soorry it will be 9C4

OpenStudy (anonymous):

this 9 choose 4 written as \[\left(\begin{matrix}9 \\ 4\end{matrix}\right)\] which is \[9!/(4!(9-4))\] solve from there. This comes from binomial theorem but this is the general multiplicative form of is \[\left(\begin{matrix}n \\ k\end{matrix}\right)\] read as n choose k and the mulitplicative form is \[n!/(k!(n-k))\]

OpenStudy (anonymous):

@azolotor ..do we have to take \[P_{4}^{9}\] too?

OpenStudy (anonymous):

I do not believe so